A bomb of mass 16 kg at rest explodes into two pieces of masses of 4 kg and 12 kg. The velocity of the 12 kg mass is 4 ms-1. The kinetic energy of the other mass is ?

Grade:11

2 Answers

Mayank
7 years ago
Pi = Pf (conservation of linear momentum)
16×0 = 12×4 + 4×v
V = 12 m/s
K.E = 1/2 m v^2
        = 1/2 × 4 × 12^2
          = 288J
Mayank pathak
7 years ago
Since initially the bomb was at rest so its momentum(mv)= 0 and no external force is applied so we can use momentum conservationLet speed of 4kg particle is v m/s0= 12×4 + 4×vV=12 m/sSo its kinetic energy= 1÷2 ×4×12= 24 J